Assume the quantity demanded for a particular commodity is given by the formula (Q)= 8000P^(-1.5). Compute the elasticity of demand

1
Expert's answer
2021-06-14T13:17:30-0400

Solution:

Elasticity of demand (Ed) ="\frac{\%\triangle Q}{\%\triangle P} = \frac{\triangle Q}{\triangle P}\times \frac{ P}{Q}"


This is a constant elasticity of demand:

Ed = "\frac{\triangle Q}{\triangle P}\times \frac{ P}{Q}"


"\frac{\triangle Q}{\triangle P} = -12,000P^{-2.5}"


"\frac{ P}{Q} = \frac{ P}{8,000P^{-1.5} }"


Ed = "-12,000P^{-2.5} \times \frac{ P}{8,000P^{-1.5} } = -1.5P^{-2.5} \times \frac{ P}{P^{-1.5} }"


= "= -1.5P^{-2.5} P P^{1.5} = -1.5P^{-2.5} P^{2.5} = -1.5"


Elasticity of demand (Ed) = -1.5